OpenStudy (anonymous):

Nine pounds of sweet potatoes cost the same as 6 pounds of apples. One pound of sweet potatoes costs twice as much as a pound of onions, while a pound of apples costs 24 cents more than a pound of onions. Find the cost of one pound of each.

OpenStudy (mathstudent55):

Let S = price per pound of sweet potatoes A = price per pound of apples N = price per pound of onions 1. "Nine pounds of sweet potatoes cost the same as 6 pounds of apples." 9S = 6A 2. "One pound of sweet potatoes costs twice as much as a pound of onions," S = 2N 3. "while a pound of apples costs 24 cents more than a pound of onions." A = N + 0.24 6A - 9S = 0 2N - S = 0 A - N = 0.24 Solve the system of equations.